# Phase reduction explains chimera shape: when multi-body interaction matters

### v1

@@ Abstract
We present an extension of the Kuramoto-Sakaguchi model for networks, deriving the second-order phase approximation for a paradigmatic model of oscillatory networks - an ensemble of non-identical Stuart-Landau oscillators coupled pairwisely via an arbitrary adjacency matrix. We explicitly demonstrate how this matrix translates into the coupling structure in the phase equations. To illustrate the power of our approach and the crucial importance of high-order phase reduction, we tackle a trendy setup of non-locally coupled oscillators exhibiting a chimera state. We reveal that our second-order phase model reproduces the dependence of the chimera shape on the coupling strength that is not captured by the typically used first-order Kuramoto-like model. Our derivation contributes to the rapidly developing field of hypernetworks, establishing a relation between the adjacency matrix and multi-body interaction terms in the high-order phase model.
@@ I Introduction
A central problem of network science is an account of multi-body interactions [1, *battiston2021, *battiston2022, *boccaletti2023], which can be responsible for various dynamical phenomena [5, *bick2023c]. In oscillatory networks, such interactions naturally appear on the level of phase description via the terms depending on the phases of three or more oscillators. In particular, such terms emerge due to the high-order phase reduction of pairwisely coupled oscillators. In this Paper, we elaborate on this scenario and treat a system of $N$ Stuart-Landau (SL) oscillators:
@@ II Main results
where the exact relation between the matrix ${\bf L}$ and coupling functions $\mathcal{R},\mathcal{S}$ is given by Eqs. (25,26) below. Recall that chimera state is a dynamical pattern with self-organized domains of synchronized (coherent) and desynchronized (incoherent) oscillators. We emphasize that in the first-order approximation, the spatial distribution of coherent and incoherent oscillators on the ring is independent of coupling strength. Indeed, if one neglects the terms $\sim\varepsilon^{2}$, goes to the co-rotating frame, and then rescales the time, $t\mapsto t/\varepsilon$, then the resulting equation does not depend on $\varepsilon$. We demonstrate that, however, the full SL model exhibits dependence of the chimera shape on $\varepsilon$, and so does the second-order phase model (3). For this, we revisit the classical model where chimera states were discovered by Kuramoto and Battogtokh [14]. Following [21], we fix the parameters of SL oscillators as $\eta_{i}=\beta_{i}=1$ and $\nu_{i}=0$. Respectively, we choose the nonlocal coupling in the form
@@
$G_{\mathrm{exp}}(x)=e^{-2x}/(1-e^{-2})\;,$ (5)
and $d_{ij}=\min\{|i-j|,N-|i-j|\}$ is the distance between the $i$th and $j$th oscillators on the ring of length $N$. Finally, we fix $\alpha_{ij}=\alpha=-\arctan(0.9)$ for all $i$ and $j$. Next, we compare the results of the numerical simulation for the SL ring with those for the first- and second-order phase models  [^2] To calculate the solution of the system (1), we used the standard fourth-order Runge-Kutta integrator with a fixed time step $dt=0.002/\varepsilon$. We discarded transients of length $2000/\varepsilon$ time units and used the next $500/\varepsilon$ time units to calculate the time-averaged phase velocities $\Omega_{i}$. The chimera state for $\varepsilon=0.01$ was obtained using the coherent-incoherent initial condition: $z_{i}=1$ for $i<0.2N$, and $z_{i}=e^{i\xi_{i}}$ otherwise, where $\xi_{i}$ is a random variable uniformly distributed in the interval $(-\pi,\pi]$. The parameter sweep in Figure 1 was calculated by adiabatically increasing $\varepsilon$ with a step of $d\varepsilon=0.01$. A similar numerical protocol was also used to calculate solutions of phase-reduced models.. The computations for the SL model clearly demonstrate that the shape of the chimera state changes for different values of $\varepsilon$ even if $\varepsilon$ remains relatively small, see Fig. 1. Moreover, the same simulations show that as $\varepsilon$ increases, the chimera states cease to exist at a fold bifurcation, while the first-order phase model does not signal this. However, if, along with the $\varepsilon^{1}$-terms, we also keep the $\varepsilon^{2}$-terms in Eq. (3), the performance of the phase reduction improves significantly. In particular, Fig. 1 shows that the second-order phase model adequately reproduces the changes in the shape of the chimera state for the variation of $\varepsilon$, which the first-order phase model cannot capture.
@@ III Second-order phase reduction for the SL network
Our derivation of Eq. (2) follows [23, 24] and consists of two main steps. The first step is the (invertible) transformation of each oscillator to its phase-isostable normal form [25, 26], expressing its state by phase $\phi_{i}\in[0,2\pi)$ and isostable amplitude $r_{i}\in\mathbb{R}$ that is zero at the limit cycle. The actual reduction of the coupled system relies on the assumption of all isostable amplitudes being enslaved by the phases, i.e., they are functions of the phases $r_{i}=R_{i}(\vec{\phi},\varepsilon)$ but not independent dynamical variables. In the second step, we suppose that the coupling is weak so that
$R_{i}(\vec{\phi},\varepsilon)=R_{i;1}(\vec{\phi})\varepsilon+\mathcal{O}(\varepsilon^{2})\,,$ (6)
and use this assumption to obtain the first and second-order terms in (2) from the formal asymptotic expansion.
For convenience, we first switch to physically meaningful parameters for each unit: frequency of the limit-cycle oscillation, $\omega_{i}=\nu_{i}-\beta_{i}$, Floquet exponent of the cycle, $\kappa_{i}=-2\eta_{i}<0$, and the non-isochronicity $\gamma_{i}=\arctan(\beta_{i})$ describing the inclination angle of isochrones at the limit cycle. With $c_{i}=\kappa_{i}\tan(\gamma_{i})/2$, Eq. (1) becomes
@@
Next, for each SL unit, we exploit the analytical formulas for phase-isostable transformation [26]:
@@
$r$ $=P(\rho)=(\rho^{2}-1)/(2\rho^{2})\,.$ (10)
We remark, that on the limit cycle ($\rho=1$) we have $r=0$, $\phi=\theta$, and $\partial_{\rho}r=\partial_{r}\rho=1$. For an autonomous system, the normal form obeys simple linear dynamics $\dot{\phi}=\omega$ and $\dot{r}=\kappa r$, is unique up to a phase shift in $\phi$ and a scaling factor in $r$, and valid in the entire basin of attraction of the limit cycle [25].
For the coupled system, using $\dot{\phi}_{i}=\partial_{\theta_{i}}\Phi_{i}\cdot\dot{\theta}_{i}+\partial_{\rho_{i}}\Phi_{i}\cdot\dot{\rho}_{i}$, $\dot{r}_{i}=\partial_{\rho_{i}}P_{i}\cdot\dot{\rho_{i}}$, and Eqs. (7-10), we write:
@@ III.1 A particular case: global coupling
Note that the frequency differences, encoded in $\Delta_{ij}=\arctan((\omega_{i}-\omega_{j})/\kappa)$, influence both the amplitude and the phase shift of the second-order coupling terms. For identical frequencies, $\Delta_{ij}=0$ and our model reduces to that of León and Pazó (up to a choice of parameters)  [^4] Using a different approach, León and Pazó derived second and third-order phase models for globally coupled identical SL oscillators [36].
@@ III.2 A particular case: non-locally coupled identical units
Now, we use formulas (14,21-23) to write an explicit form of the second-order phase approximation for a ring of identical SL oscillators, i.e., for the chimera setup.
We take $\omega_{i}=\omega$, $\gamma_{i}=\gamma$, and $\kappa_{i}=\kappa$  [^5] We remind that these parameters are related to the parameters $\eta_{i}$, $\beta_{i}$ and $\nu_{i}$ in Eq. (1) by formulas $\omega_{i}=\nu_{i}-\beta_{i}$, $\kappa_{i}=-2\eta_{i}$ and $\gamma_{i}=\arctan(\beta_{i})$.. Then, $\Delta_{ij}=0$. The non-local coupling is organized in the standard way:
@@ IV Discussion
Finally, we mention that general results [24] provide a route to computing third- and higher-order terms, leading to quadruplet and so on terms, though this task is tedious and highly time-consuming. Another conclusion is that multi-body interaction appears for networks of arbitrary limit-cycle oscillators, though the closed analytical derivation of the second-order terms becomes unfeasible and must be partially done numerically, cf. [24]. We emphasize that there exists an entirely different scenario resulting in multi-body interaction terms due to nonlinear coupling [31, *Komarov-Pikovsky-15a, *ashwin2016, *Pikovsky-Rosenblum-22] which we do not address here.

### v2

@@ Abstract
We present an extension of the Kuramoto-Sakaguchi model for networks, deriving the second-order phase approximation for a paradigmatic model of oscillatory networks - an ensemble of non-identical Stuart-Landau oscillators coupled pairwisely via an arbitrary adjacency matrix. We explicitly demonstrate how this matrix translates into the coupling structure in the phase equations. To illustrate the power of our approach and the crucial importance of high-order phase reduction, we tackle a trendy setup of non-locally coupled oscillators exhibiting a chimera state. We reveal that our second-order phase model reproduces the dependence of the chimera shape on the coupling strength that is not captured by the typically used first-order Kuramoto-like model. Our derivation contributes to a better understanding of complex networks’ dynamics, establishing a relation between the adjacency matrix and multi-body interaction terms in the high-order phase model.
@@ I Introduction
A central problem of network science is an account of multi-body interactions [1, *battiston2021, *battiston2022, *boccaletti2023], which can be responsible for various dynamical phenomena [5, *bick2023c]. In oscillatory networks, such interactions naturally appear on the level of phase description via the terms depending on the phases of three or more oscillators. In particular, such terms emerge due to the high-order phase reduction of pairwisely coupled oscillators. In this Letter, we elaborate on this scenario and treat a system of $N$ Stuart-Landau (SL) oscillators:
@@ II Main results
where the exact relation between the matrix ${\bf L}$ and coupling functions $\mathcal{R},\mathcal{S}$ is given by Eqs. (25,26) below. Recall that chimera state is a dynamical pattern with self-organized domains of synchronized (coherent) and desynchronized (incoherent) oscillators. We emphasize that in the first-order approximation, the spatial distribution of coherent and incoherent oscillators on the ring is independent of coupling strength. Indeed, if one neglects the terms $\sim\varepsilon^{2}$, goes to the co-rotating frame, and then rescales the time, $t\mapsto t/\varepsilon$, then the resulting equation does not depend on $\varepsilon$. We demonstrate that, however, the full SL model exhibits dependence of the chimera shape on $\varepsilon$, and so does the second-order phase model (3). For this, we revisit the classical model where chimera states were discovered by Kuramoto and Battogtokh [14]. We choose the nonlocal coupling in the form
@@
$G_{\mathrm{exp}}(x)=e^{-2x}/(1-e^{-2})$ (5)
is an exponentially decaying function associated with the adiabatic elimination of the diffusive mediator in a two-component system, and $d_{ij}=\min\{|i-j|,N-|i-j|\}$ is the distance between the $i$th and $j$th oscillators on the ring of length $N$. (Note the normalization factors in (4) and (5) guarantee that the interaction is asymptotically balanced, i.e., $\sum_{j=1}^{N}L_{ij}=0$ for $N\to\infty$.) As for other parameters, we adopt them from [21]. Namely, we fix the parameters of SL oscillators as $\eta_{i}=\beta_{i}=1$ and $\nu_{i}=0$, and we choose $\alpha_{ij}=\alpha=-\arctan(0.9)$ for all $i$ and $j$. Next, we compare the results of the numerical simulation for the SL ring with those for the first- and second-order phase models  [^2] To calculate the solution of the system (1), we used the standard fourth-order Runge-Kutta integrator with a fixed time step $dt=0.002/\varepsilon$. We discarded transients of length $2000/\varepsilon$ time units and used the next $500/\varepsilon$ time units to calculate the time-averaged phase velocities $\Omega_{i}$. The chimera state for $\varepsilon=0.01$ was obtained using the coherent-incoherent initial condition: $z_{i}=1$ for $i<0.2N$, and $z_{i}=e^{i\xi_{i}}$ otherwise, where $\xi_{i}$ is a random variable uniformly distributed in the interval $(-\pi,\pi]$. The parameter sweep in Figure 1 was calculated by adiabatically increasing $\varepsilon$ with a step of $d\varepsilon=0.01$. A similar numerical protocol was also used to calculate solutions of phase-reduced models.. The computations for the SL model clearly demonstrate that the shape of the chimera state changes for different values of $\varepsilon$ even if $\varepsilon$ remains relatively small, see Fig. 1. Moreover, the same simulations show that as $\varepsilon$ increases, the chimera states cease to exist at a fold bifurcation, while the first-order phase model does not signal this. However, if, along with the $\varepsilon^{1}$-terms, we also keep the $\varepsilon^{2}$-terms in Eq. (3), the performance of the phase reduction improves significantly. In particular, Fig. 1 shows that the second-order phase model adequately reproduces the changes in the shape of the chimera state for the variation of $\varepsilon$, which the first-order phase model cannot capture.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ III Second-order phase reduction for the SL network
Our derivation of Eq. (2) follows [23, 24] and consists of two main steps. The first step is the (invertible) transformation of each oscillator to its phase-isostable normal form [25, 26], expressing its state by phase $\phi_{i}\in[0,2\pi)$ and isostable amplitude $r_{i}\in\mathbb{R}$ that is zero at the limit cycle. The actual phase reduction relies on the weak coupling assumption. In this case, our coupled SL system with $\eta_{i}>0$ has an attractive smooth invariant $N$-dimensional torus [27] parameterized by the phases $\phi_{i}$. Thus, focusing on the system’s long-term behavior, we take all isostable amplitudes as functions of the phases, i.e., $r_{i}=R_{i}(\vec{\phi},\varepsilon)$. In the second step, we suppose that each $R_{i}(\vec{\phi},\varepsilon)$ is smooth and $R_{i}(\vec{\phi},0)=0$ so that
$R_{i}(\vec{\phi},\varepsilon)=R_{i;1}(\vec{\phi})\varepsilon+\mathcal{O}(\varepsilon^{2})\,.$ (6)
Then, we use ansatz (6) to obtain the first and second-order terms in (2) from the formal asymptotic expansion.
For convenience, we first switch to physically meaningful parameters for each unit: frequency of the limit-cycle oscillation, $\omega_{i}=\nu_{i}-\eta_{i}\beta_{i}$, Floquet exponent of the cycle, $\kappa_{i}=-2\eta_{i}<0$, and the non-isochronicity parameter $\gamma_{i}=\arctan(\beta_{i})$.
With $c_{i}=\kappa_{i}\tan(\gamma_{i})/2$, Eq. (1) becomes
@@
It is well-known [26] that for each SL oscillator of the form $\dot{\theta}=\omega+c(\rho^{2}-1)$, $\dot{\rho}=\kappa\rho(\rho^{2}-1)/2$, there exists a phase-isostable transformation:
@@
$r$ $=P(\rho)=(\rho^{2}-1)/(2\rho^{2})\,$ (10)
with $\tan(\gamma)=2c/\kappa$, which is defined by two characteristic properties. First, in the new coordinates, the dynamics of the amplitude $r$ and the phase $\phi$ decouple and take on a simple linear form $\dot{\phi}=\omega$ and $\dot{r}=\kappa r$. Second, on the limit cycle ($\rho=1$), it holds $r=0$ and $\phi=\theta$. Note that the transformation (9,10) defined in this way  [^3] The phase-isostable transformation (9,10) is obtained by solving two differential equations $\dot{\phi}=\omega=\frac{d}{dt}\Phi(\theta,\rho)$ and $\dot{r}=\kappa r=\frac{d}{dt}P(\rho)$, which explicitly read $\omega=(\omega+c(\rho^{2}-1))\partial_{\theta}\Phi+0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}\Phi$ and $\kappa P=0.5\kappa\rho(\rho^{2}-1)\partial_{\rho}P$, with two boundary conditions $\Phi(\theta,1)=\phi$ and $P(1)=0$. Note that although Eq. (9) is a unique solution, Eq. (10) is defined only up to a scaling factor. is isomorphic on the entire basin of attraction of the limit cycle, i.e. for all $\rho>0$ [25]. In addition, the isochrones of the SL oscillator $\Phi(\theta,\rho)=\mathrm{const}$ are logarithmic spirals with the inclination angle $\gamma$ relative to the outward normal on the limit cycle ($\rho=1$). For $\gamma=0$, these spirals turn into straight rays and Eq. (9) degenerates into the identity $\phi=\theta$.
Using (9,10) and the identities $\dot{\phi}_{i}=\partial_{\theta_{i}}\Phi_{i}\cdot\dot{\theta}_{i}+\partial_{\rho_{i}}\Phi_{i}\cdot\dot{\rho}_{i}$ and $\dot{r}_{i}=\partial_{\rho_{i}}P_{i}\cdot\dot{\rho_{i}}$, we write system (7,8) in the form:
@@ III.1 A particular case: global coupling
Note that the frequency differences, encoded in $\Delta_{ij}=\arctan((\omega_{i}-\omega_{j})/\kappa)$, influence both the amplitude and the phase shift of the second-order coupling terms. For identical frequencies, $\Delta_{ij}=0$, our model reduces to that of León and Pazó (up to a choice of parameters)  [^5] Using a different approach, León and Pazó derived second and third-order phase models for globally coupled identical SL oscillators [38].
@@ III.2 A particular case: non-locally coupled identical units
Now, we use formulas (14,21-23) to write an explicit form of the second-order phase approximation for a ring of identical SL oscillators, i.e., for the chimera setup. We take $\omega_{i}=\omega$, $\gamma_{i}=\gamma$, and $\kappa_{i}=\kappa$  [^6] We remind that these parameters are related to the parameters $\eta_{i}$, $\beta_{i}$ and $\nu_{i}$ in Eq. (1) by formulas $\omega_{i}=\nu_{i}-\eta_{i}\beta_{i}$, $\kappa_{i}=-2\eta_{i}$ and $\gamma_{i}=\arctan(\beta_{i})$.. Then, $\Delta_{ij}=0$. The non-local coupling is organized in the standard way:
@@ IV Discussion
Finally, we mention that general results [24] provide a route to computing third- and higher-order terms, leading to quadruplet and so on terms, though this task is tedious and highly time-consuming. Another conclusion is that multi-body interaction appears for networks of arbitrary limit-cycle oscillators, though the closed analytical derivation of the second-order terms becomes unfeasible and must be partially done numerically, cf. [24]. We emphasize that there exists an entirely different scenario resulting in multi-body interaction terms due to nonlinear coupling [33, *Komarov-Pikovsky-15a, *ashwin2016, *Pikovsky-Rosenblum-22] which we do not address here.